OpenStudy (anonymous):

PLZZZ HELP ILL FAN AND ILL TRY TO MEDAL Transform the following equation written in standard form into slope-intercept form: 2x+3y=5. a) y=(32/)x+5 b) y=−(23/)x+5 c) y=−(23/)x+53/ d) y=−2x+5 Given the following equation: −3x+4y=5. This line has a positive slope and intersects the y-axis at (0,54/) . Question 4 options: True False What is the slope of the line that passes through the points A (2,3) and B (-1,-3)? (Your answer will be a number.)

rebeccaxhawaii (rebeccaxhawaii):

slope intercept form is y=mx+b

OpenStudy (anonymous):

@likeabossssssss

likeabossssssss (likeabossssssss):

what do u yhink it is

likeabossssssss (likeabossssssss):

think***

OpenStudy (anonymous):

uhh well #1 I think is either d or c? and #2 I pretty sure it false and #3 I dunno

OpenStudy (anonymous):

slope intercept form is y=mx+b like @rebeccaxhawaii said earlier so which one looks like that the most?

OpenStudy (anonymous):

D

rebeccaxhawaii (rebeccaxhawaii):

yes that is correct

OpenStudy (anonymous):

^

OpenStudy (anonymous):

wait a or d ?

rebeccaxhawaii (rebeccaxhawaii):

wait no it isnt

OpenStudy (anonymous):

You can double check by pretending the inequality is an equal sign and isolating the variable "y"

rebeccaxhawaii (rebeccaxhawaii):

because 3y we cant just get rid of the 3 can you send a screenshot of the question

rebeccaxhawaii (rebeccaxhawaii):

and choices

OpenStudy (anonymous):

I cant but ill re write it because I see I mess up on the anwers

OpenStudy (anonymous):

Transform the following equation written in standard form into slope-intercept form: 2x+3y=5. Question 3 options: a) y=(3/2)x+5 b) y=−(2/3)x+5 c) y=−(2/3)x+53/ d) y=−2x+5

OpenStudy (anonymous):

C)y=-(2/3)x+5/3

OpenStudy (anonymous):